Suppose the demand for commodity X as the function of its own price and consumer’s money income,

keeping other factors constant is described by the following equation:

X = 100,000 – 2PX + 7M; where M is consumer’s money income in Birr and PX is price of the

commodity X in Birr. If the price is Birr 10,000 and money income is Birr 15,000, then:

a. what is the total demand for X at the prevailing market price and income

b. calculate the income elasticity of demand for the commodity &; comment on the type of elasticity and

type of commodity

c. calculate the price elasticity of demand for the commodity and determine the type of elasticity


1
Expert's answer
2022-01-25T08:13:34-0500

"a) X=100000- 2PX+7M"


PX=10000


"M=15000"

"X=100,000-2(10,000)+7(15,000)" "X=100,000-20,000+105,000"

"X=80,000+105,000=185,000"

b) Income elasticity

"\frac{∆X}{∆I} × \frac{I}{X}"

"\frac{∆X}{∆I} =7",(change in money income given in the function)

"I=15000\nX=185000"

"7\u00d715000\/185000=0.567"

A 1% increase in income results in the quantity demanded of good X increasing by 0.567. The elasticity is unit elastic since it is approximately equal to 1 and the commodity is a normal good since the elasticity maintains a positive value.

c) Price elasticity

"\frac{∆Q}{∆P} ×\frac{P}{Q}"

"\frac{∆Q}{∆P}=-2" (change in price from the function given)

"P=10000"

"-2\u00d710000\/185000=-0.108"

Since the value is less than 1, it is inelastic.